# Ink and Individuality: Crafting a Personalised Narrative in the Age of LLMs

### v1

@@ Abstract.
Individuality and personalization comprise the distinctive characteristics that make each writer unique and influence their words in order to effectively engage readers while conveying authenticity. However, our growing reliance on LLM-based writing assistants risks compromising our creativity and individuality over time. We often overlook the negative impacts of this trend on our creativity and uniqueness, despite the possible consequences. This study investigates these concerns by performing a brief survey to explore different perspectives and concepts, as well as trying to understand people’s viewpoints, in conjunction with past studies in the area. Addressing these issues is essential for improving human-computer interaction systems and enhancing writing assistants for personalization and individuality.
copyright: none
@@ 1. Introduction
The exceptional performance of large language models (LLMs) across a wide range of applications makes them increasingly well-liked in academia and industry (Chang et al., 2023). Especially with chatbots such as Microsoft Copilot, Gemini, and ChatGPT, which have the transformative capability of producing polished and well-crafted stories, essays and assignments akin to those written by a professional novelist or a writer (Chung et al., 2022; Yuan et al., 2022). Not to mention, other writing tools such as Quillbot and Grammarly have made writing a straightforward, time-efficient task now. With recent works such as (Kim et al., 2023) and (Yeh et al., 2024), we are at a junction where personalization is of high significance with LLM usage for writing. This raises a question: Are users concerned about the level of personalization in their writings with the emergence of AI-assisted writing tools? In this paper, we explore (i) the importance of personalization and individuality of writing and (ii) personalization in the age of LLMs. In this study of these issues, we try to understand the perspective of individuals through a thoroughly conducted survey, along with previous research work done in this domain. We aim to contribute to the development of personalization in writing with LLMs along with its significance and try to highlight whether individuals are affected by it. It is important to highlight the personal touch of an individual when writing content, for example, crafting personal statement essays for university applications or cover letters for jobs, which are important factors in whether someone gets accepted or rejected.
@@ 2. Personalization and Individuality in Writing
3. Personalization in the Age of LLMs
@@ 4. Problems and Perspectives
To investigate the role of LLMs in content personalization and individuality, we conducted a brief survey with 35 participants. The demographic profile of the participants are available in Section A. The responses that we received show different perspectives and concerns in this area. Figure 1 shows the frequency of LLM assistance in three types of writing: personal writing (letters, messages and birthday wishes), formal writing (CVs and applications), and academic writing (reports, assignments and homework).
The term ”Personal Touch” in writing, in this paper, refers to the unique voice, style, and perspective that an individual writer brings to their work. It encompasses a distinct tone, relatable personal anecdotes or experiences, clear opinions and viewpoints, preferred writing style and word choice, as well as the ability to forge an emotional connection with the reader. The personal touch is what allows a writer’s personality and humanity to shine through their words, distinguishing their writing from others and making it feel more authentic, intimate, and engaging.
@@
So, do LLMs negatively affect personalization and individuality? Our survey data reveals that a majority (71.4%) value personal touches in writing (Figure 2 (a)), and many believe LLMs can capture their personal touch (Figure 2 (b)). However, despite this, 60% of respondents feel their personal touch is lost when using LLMs (Figure 2 (c)), suggesting a discrepancy between perception and reality regarding the impact of LLMs on personalization and individuality.
@@
Although participants acknowledge the significance of personal touches in writing and think LLMs can capture these touches, a significant number of the respondents still feel that their personal touch is lost when using LLMs. This discrepancy raises questions about possible limitations in current LLMs’ ability to accurately replicate individual style and expression. One of the reasons for believing LLMs can capture personal touches is similarity in tone or structure. However, the participants actually feel their personal touch is lost because nuances in expression or unique writing quirks are not faithfully replicated, leading to a sense of impersonal or generic output. Another possible explanation is that participants perceive a decrease in their writing or thinking ability as they put less effort, leading them to feel concerned that their personal touch is being lost.
Furthermore, these gaps emphasize the need for more research and development to address the challenges associated with maintaining personalization and individuality in the age of AI writing assistants. It also raises questions about the effectiveness of LLMs in preserving the authenticity and uniqueness of personal writing, suggesting that reliance on these tools may lead to a perceived loss of personalization and individuality despite attempts to integrate personal touches.
@@ 5. Why does losing our individuality matter?
6. Discussions on Limitations and Future Works
We acknowledge that there are some shortcomings in the survey questionnaire regarding quantifying the definitions of options and word choices for certain questions. As this is an ongoing endeavor, we aim to rectify these issues in future surveys and data collections. We also acknowledge that exploring the correlation between perception towards the need and loss of personal touch with each of these domains would be insightful. We will certainly consider incorporating this aspect into future analyses. Further, we notice a gap regarding the lack of clarity on what each participant perceives as a ’personal touch.’ Indeed, understanding individual perspectives on this matter could significantly enhance the depth and context of our analysis, particularly in relation to questions concerning the capability of LLMs to capture personal touch and the perceived loss of personal touch when using LLMs. We will clarify these issues to the further survey participants.
@@ 6. Discussions on Limitations and Future Works
This study investigates the different obstacles and perspectives of writing with personalization and individuality in the age of LLMs. According to our survey, we are currently heavily reliant on LLMs for writing any kind of content, from personal to format to academic; and this heavy use of LLMs is harming our individuality, but we are not too concerned about it. This study connects our survey with literature to understand individual perspectives and contribute to the development of personalised writing with LLMs.
Below is a concise summary of both the data-supported findings regarding participants’ motives and some speculative insights:
  • We investigate whether writers are concerned about adding personalized content to their writings and if LLMs can capture it. We find out most writers have claimed that personal touch is important in writing, but LLMs fail to capture their persona and nature for content writing.
@@
  • We emphasize the need for features or modes in current LLMs that will help obtain the personal touch in writing to embrace an individual’s uniqueness. Future discussions and research on this area can focus on this aspect.

### v5

@@ Abstract.
Individuality and personalization comprise the distinctive characteristics that make each writer unique and influence their words in order to effectively engage readers while conveying authenticity. However, our growing reliance on LLM-based writing assistants risks compromising our creativity and individuality over time. We often overlook the negative impacts of this trend on our creativity and uniqueness, despite the possible consequences. This study investigates these concerns by performing a brief survey to explore different perspectives and concepts, as well as trying to understand people’s viewpoints, in conjunction with past studies in the area. Addressing these issues is essential for improving human-centered aspects of AI models and enhancing writing assistants for personalization and individuality.
Human computer interaction, Writing Assistants, Personalization and Individuality, Creativity and Cognition, Large Language Models
journalyear: 2024copyright: acmlicensedconference: The Third Workshop on Intelligent and Interactive Writing Assistants; May 11, 2024; Honolulu, HI, USAbooktitle: The Third Workshop on Intelligent and Interactive Writing Assistants (In2Writing ’24), May 11, 2024, Honolulu, HI, USAdoi: 10.1145/3690712.3690724isbn: 979-8-4007-1031-5/24/05ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Language resourcesccs: Computing methodologies Discourse, dialogue and pragmatics
@@ 1. Introduction
The exceptional performance of large language models (LLMs) across a wide range of applications makes them increasingly well-liked in academia and industry (Chang et al., 2023). Especially with chatbots such as Microsoft Copilot, Gemini, and ChatGPT, which have the transformative capability of producing polished and well-crafted stories, essays and assignments akin to those written by a professional novelist or a writer (Chung et al., 2022; Yuan et al., 2022). Not to mention, other writing tools such as Quillbot and Grammarly have made writing a straightforward, time-efficient task now. With recent works such as (Kim et al., 2023) and (Yeh et al., 2024), we are at a junction where personalization is of high significance with LLM usage for writing. This raises a question: Are users concerned about the level of personalization in their writings with the emergence of AI-assisted writing tools? (Figure 1)
Figure 1. Crafting a Personalised Narrative in the Age of LLMs: Is it better or worse?
In this paper, we explore (i) the importance of personalization and individuality of writing and (ii) personalization in the age of LLMs. In this study of these issues, we try to understand the perspective of individuals through a thoroughly conducted survey, along with previous research work done in this domain. We aim to contribute to the development of personalization in writing with LLMs along with its significance and try to highlight whether individuals are affected by it. It is important to highlight the personal touch of an individual when writing content, for example, crafting personal statement essays for university applications or cover letters for jobs, which are important factors in whether someone gets accepted or rejected.
@@ 2. PERSONALIZATION AND INDIVIDUALITY IN WRITING
In this work, we use the term ”Personal Touch” frequently. This term refers to the unique voice, style, and perspective that an individual writer brings to their work. It encompasses a distinct tone, relatable personal anecdotes or experiences, clear opinions and viewpoints, preferred writing style and word choice, as well as the ability to forge an emotional connection with the reader. The personal touch is what allows a writer’s personality and humanity to shine through their words, distinguishing their writing from others and making it feel more authentic, intimate, and engaging.
3. Personalization in the Age of LLMs
@@
Figure 4. Average prompt length by survey participants
@@ 4. Problems and Perspectives
4.1. Survey Information
To investigate the role of LLMs in content personalization and individuality, we conducted a brief survey with 35 participants. The demographic profile of the participants are available in Appendix A. The responses that we received show different perspectives and concerns in this area. Figure 2 shows the frequency of LLM assistance in three types of writing: personal writing (letters, messages and birthday wishes), formal writing (CVs and applications), and academic writing (reports, assignments and homework).
4.2. Survey Questions
Below are the survey questions included in the study:
  • Personal Writing: How often do you write personal notes (letters, messages, birthday wishes) using ChatGPT/other AI writing assistants?
    Options: Every time, Most of the time, Often, Some times, Never.
  • Formal Writing: How often do you use it for formal writings (job tasks, applications and stuffs)?
    Options: Every time, Most of the time, Often, Some times, Never.
  • Academic Writing: How often do you use it for academics (reports, assignments, anything)?
    Options: Every time, Most of the time, Often, Some times, Never.
  • Average prompt length by survey participants : Your average prompt length (word count) (to assess context and modifications)
    Options: less than 10 words, 10-20 words, 20-50 words, 50+ words
  • Can LLMs catch you personal touch? : Do you think AI assistants can catch your personal touch (your writing style, or feelings)? (As they generate general outputs most of the times)
    Options: Every time, Most of the time, Often, Some times, Never.
  • Does personal touch matter? Do you think your personal touch in writing matters in different types of writing?
    Options: Yes, No, sometimes.
  • Is your personal touch lost when using LLMs? Do you think your personal touch in writing is being lost while using AI?
    Options: Yes, No, sometimes.
4.3. Survey Statistics and Interpretation
@@
So, do LLMs negatively affect personalization and individuality? Our survey data reveals that a majority (71.4%) value personal touches in writing (Figure 3 (a)), and many believe LLMs can capture their personal touch (Figure 3 (b)). However, despite this, 60% of respondents feel their personal touch is lost when using LLMs (Figure 3 (c)), suggesting a discrepancy between perception and reality regarding the impact of LLMs on personalization and individuality. Also, figure 4 illustrates that the majority of prompts are relatively short in length, suggesting that participants may not have employed sufficient prompt engineering techniques to generate more personalized or engaging content.
@@
4.4. Observation
Respondents recognize the value of personalization in writing and believe LLMs can capture some aspects of personal style, yet many still feel that their unique touch gets lost when using LLMs. This discrepancy highlights potential limitations in current LLMs’ ability to faithfully replicate individual expression. One of the reasons for believing LLMs can capture personal touches is similarity in tone or structure. However, participants feel that their personal style is lost because the individuality in their expression or unique writing habits are not accurately reflected, resulting in a sense of impersonal or generic writing for all. Additionally, participants may feel a decline in their own writing or thinking skills when relying on LLMs, contributing to concerns about losing personal touch. These findings underscore the need for more research and development to address challenges in maintaining personalization in AI writing tools and raise doubts about LLMs’ effectiveness in preserving authenticity and uniqueness in personal writing.
@@ 5. Why does losing our individuality matter?
6. Related Works and Discussion
From our surveys, we have established the need for LLMs to capture personalization and reflect in for the various forms of writing: personal, formal and academic. With most of our survey indicating users want personified writing, we hope LLMs which aides in such personalization are constructed because with LLMs becoming a part of our daily and professional lives, it is imperative these AI-assisted writing tools are not giving us monotonous and generic content. One such collaborative group-AI Brainwriting brainstorming framework was developed by Shaer et al. (2024). They included an LLM to improve the group ideation process, and then they assessed the process of idea generation along with the resulting solution space. With creativity being a main priority to be included with LLMs, Xu et al. (2024) propose Jamplate, a digital whiteboard plugin with LLM integration, to simplify the idea generating process which is an an alternate strategy that substitutes chat interfaces with structured templates is put forth. In their preliminary research, they discovered Jamplate helps inexperienced designers think more critically and improve their ideas. With such works existing to tackle the existing issues of LLMs, we believe our work is a step towards the right direction and brings attention to address the lack of creativity and human-touch like feel to automated content writing. Additionally, this is a crucial topic for discussion considering the presence of other widely spoken languages globally such as Chinese, French, German, Hindi, and more. Given the widespread international use of LLMs, this issue requires prompt attention and resolution.
Limitations. One drawback of this study is the limited sample size of only 35 participants. Enhancing the number of participants could enhance the robustness and significance of the research outcomes. Also, the survey questions can be more effectively designed. We hope to work on these issues in future.
@@ 7. Conclusion
This study investigates the different obstacles and perspectives of writing with personalization and individuality in the age of LLMs. According to our survey, we are currently heavily reliant on LLMs for writing any kind of content, from personal to format to academic; and this heavy use of LLMs is harming our individuality, but we are not too concerned about it. This study connects our survey with literature to understand individual perspectives and contribute to the development of personalised writing with LLMs. Below is a concise summary of both the data-supported findings regarding participants’ motives and some speculative insights:
  • We investigate whether writers are concerned about adding personalized content to their writings and if LLMs can capture it. This research suggests a discrepancy between the perceived and actual impact of LLMs on personalization and individuality in writing. We find out most writers have claimed that personal touch is important in writing, but LLMs fail to capture their persona and nature for content writing, indicating that extensive reliance on LLMs homogenizes content and diminishes individual expressions.
@@
In observation, we emphasize the need for features or modes in current LLMs that will help obtain the personal touch in writing to embrace an individual’s uniqueness. Future discussions and research on this area can focus on this aspect.
Author Contributions
ATW conceived the core idea, developed the methodology, designed the experiments and the data collection, performed some experiments, conducted all the analyses, prepared visualizations, wrote and edited core parts of the paper, and led the whole project. RI helped in background study, formal analysis, initial draft writing and editing. MRI collected the data and was involved in editing.